OpenStudy (anonymous):

The ordered pair (-4,3) is located on a circle in the coordinate plane given by the equation (x+1)^2 + (y+1)^2 = r^2. If r is the radius of the circle, which of the ff points must also lie on the circle? a. (-6,-2.8) b. (-3.9,2) c. (1,2.7) d. (-1,-4) e. (2,3)

OpenStudy (anonymous):

From the given point you can compute r

OpenStudy (anonymous):

\[(-4+1)^2 + (3+1)^2 = r^2\\ 9+16=r^2\\ r=5 \]

OpenStudy (anonymous):

It seem obvious (2,3) on the circle \[ (x+1)^2 + (y+1)^2 = 25\\ (2+1)^2 + (3+1)^2 = 25\\ 3^2 +4^2=25 \]

OpenStudy (anonymous):

thank youuu c:

OpenStudy (anonymous):

YW